40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 17 12:32:35 2021, Koji, Update, SUS, Remaining task for 2021 
    Reply  Fri Dec 17 19:19:42 2021, Koji, Update, SUS, Remaining task for 2021 PXL_20211218_030145356.MP.jpg
       Reply  Sun Dec 19 07:52:51 2021, Anchal, Update, SUS, Remaining task for 2021 
    Reply  Wed Dec 22 18:16:23 2021, Koji, Update, SUS, Remaining task for 2021 ITMX_2016_07_21.jpgScreen_Shot_2021-12-22_at_18.03.42.png
Message ID: 16534     Entry time: Wed Dec 22 18:16:23 2021     In reply to: 16519
Author: Koji 
Type: Update 
Category: SUS 
Subject: Remaining task for 2021 

The in-vacuum installation team has reported that the side OSEMs of ITMX and LO1 are going to be interfering if place LO1 at the planned location.
I confirmed that ITMX has the side magnet on the other side (Attachment 1 ITMX photo taken on 2016/7/21). So we can do this swap.

The ITMX side OSEM is sticking out most. By doing this operation, we will recover most of the space between the ITMX and LO1. (Attachment 2)

Attachment 1: ITMX_2016_07_21.jpg  209 kB  | Hide | Hide all
ITMX_2016_07_21.jpg
Attachment 2: Screen_Shot_2021-12-22_at_18.03.42.png  2.425 MB  | Hide | Hide all
Screen_Shot_2021-12-22_at_18.03.42.png
ELOG V3.1.3-